The linear density of a string is 1.5 x 10^-4 kg/m
Answer #1

The wave equation for a transverse wave is
y(x,t) = ym sin ( kx - \omega t )
Where y(x,t) is the displacement of an element of the wave
Wave speed v = \omega / k
Where \omega is the angular frequency
f = 23 s-1 , \omega  = 23 rad /s
k = 1.6 m-1
v = 23 / 1.6 = 14.375 m / s
The tension in the string
T = v2\mu
Where T is the tension in the string and \mu is the linear density of the string
T = (14.375)2 x 1.5 x 10-4
T = 0.030996 kg m /s2
or
T = 0.0310 N